OpenStudy (anonymous):

Check whether the series defined below is convergent and if so, find its sum.

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ 2^n }{ (-12)^{n-1} }\] and the equation.

OpenStudy (anonymous):

What math class is this for?

OpenStudy (anonymous):

likely calc 2, because i just had a test on this

OpenStudy (anonymous):

BTW, this series converges.

OpenStudy (anonymous):

Math 1B in Australia (presumably similar to calc 2 from what i've seen).

OpenStudy (anonymous):

ok, the sum will also be -12/5

OpenStudy (anonymous):

thanks, good to know but how do you work that out?

OpenStudy (anonymous):

Take the limit as n goes to infinity.

OpenStudy (anonymous):

thus far i've found the absolute value, then tried to find the limit as it approaches infinity. That's what i believe you do.

OpenStudy (anonymous):

Ya. That would be correct. So, the form is an+1/an

OpenStudy (anonymous):

And you plug in the formula that you have into this equation.

OpenStudy (anonymous):

-(2^n+1)/(12^(n+1)(-1)) all over (2^n)/(-12^(n-1))

OpenStudy (anonymous):

okay so because the is to a higher power the limit is equal to 0 and thus converges. Hate to ask but how do you find the sum. Think you tried to explain it there but i'm still a bit confused. does: (-(2^n+1)/(12^(n+1)(-1)))/((2^n)/(-12^(n-1))) find me the sum when n=1?